How much energy (in Joules) will be produced on complete fission of 1 kg of Uranium (235U), considering that only 0.1% of mass is converted into energy.



1
Expert's answer
2021-12-16T18:03:42-0500

The mass converted into energy is "m = 1kg\cdot 0.001 = 0.001kg". Using the energy-mass equivalance, find:


"E = mc^2"

where "c = 3\times 10^{8}m/s" is the speed of light. Thus, obtain:


"E = 0.001kg\cdot (3\times 10^8m/s)^2 = 9\times 10^{13}J"

Answer. "9\times 10^{13}J".